# ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction

### v1

@@ 3.2.2 Analysis of Preference Clarity Based the Entropy of Directed Graph Structures
where $\tau(G_{i})$ is the normalized structural entropy for the graph $G_{i}$ corresponding to question $q_{i}$, and $|\mathbf{Q}|$ is the total number of questions in the set $\mathbf{Q}$. A higher value of $\tau_{\text{avg}}$ signifies that the preference relations generated by the Evaluator LLM are more complex, indicating a lower level of overall preference clarity.
3.3 Filtering Strategy for Preference Data That Induces Non-Transitivity
After decomposing a cyclic directed graph into multiple SCCs, if the cycles within each SCC can be eliminated and each SCC can be reconstructed as a DAG, then the entire directed graph can likewise be transformed into a Directed Acyclic Graph (DAG).
Building upon this observation, our framework transforms cyclic directed graphs into DAGs by converting each SCC into a DAG, preserving inter-SCC preference relations. This process ensures the global graph structure becomes acyclic without altering inter-component preferences. Subsequently, we derive transitive preference relations from the final DAG to filter Evaluator LLMs’ preference data, facilitating self-purification of the training data. The detailed procedure is as follows:
@@ 3.3 Filtering Strategy for Preference Data That Induces Non-Transitivity
[…]
@@
    The training data is traversed sequentially, and all data consistent with the correct preferences are added to the dataset "Cleaned".
See Appendix E for detailed algorithm information.
@@ 4 Experiment Setup
Figure 3: Comparison of data volumes between "Raw" and "Cleaned" training sets across datasets. The "Cleaned" dataset’s volume is approximately 80% of the "Raw" dataset for each dataset.
@@
Qwen-Base 82.83% 0.9221 78.93% 0.8912 83.36% 0.9138 80.96% 0.9097 81.46% 0.9123
Qwen-Raw 62.02% 0.7963 57.50% 0.8034 55.78% 0.7734 64.29% 0.8159 59.69% 0.7665
Qwen-Random 60.58% 0.7899 63.57% 0.8078 58.81% 0.7711 60.35% 0.8116 57.14% 0.7644
Qwen-Cleaned (ours) 44.85% 0.6997 43.93% 0.7262 46.96% 0.6942 48.53% 0.7149 48.98% 0.6802
Table 1: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw", "Random", and "Cleaned" training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base dataset.
Dataset: In this study, we conduct experimental validation using the AlpacaEval benchmark (Li et al., 2023). AlpacaEval is specifically designed to assess the overall capabilities of LLMs in open-domain, instruction-following tasks, covering a wide range of evaluation scenarios such as reasoning and text generation. The benchmark comprises five datasets: Helpful_Base, Oasst, Koala, Vicuna, and Self-Instruct.
Teacher Model Selection: We evaluated the performance of Qwen2.5-MaxQwen Team (2025) using the 2.5k human-annotated samples provided by AlpacaEval, employing the Chain-of-Thought Comparison (CoT Comparison) Li et al. (2023) . The detailed prompt template is provided in Appendix C.1. Qwen2.5-Max demonstrated exceptional performance in both human agreement rate, achieving an agreement level of 68.9%. Based on these results, we selected Qwen2.5-Max as our teacher model.
Experiment Details: We selected 21 representative LLMs from the AlpacaEval leaderboard and divided them into two groups based on the experimental design: 14 LLMs were used to collect preference data for constructing the training set, while the remaining 7 LLMs were employed to build the testing set for performance validation(see Appendix A for details). Throughout the experiment, we consistently applied the CoT Comparison prompt template for response generation, with the decoding temperature fixed at 0. Preference data for constructing the training set was generated by invoking the official Qwen2.5-Max API. The unfiltered Raw dataset was labeled as "Raw", and subsequently filtered according to the steps described in Section 3.3, resulting in the "Cleaned" dataset. The data distribution across datasets is illustrated in Figure 3. For comparative analysis, we fine-tuned models using both the "Raw" and "Cleaned" training sets derived from five datasets, ultimately obtaining a total of 10 fine-tuned models.
In the experiments, we used Qwen2.5-7B-Instruct as the base model and fine-tuned it using the Low-Rank Adaptation (LoRA) methodHu et al. (2022), with a rank of 8. The model was trained for 3 epochs, with a learning rate of $1\times 10^{-4}$ and a batch size of 16.
@@ 5 Results and Analysis
In this section, we validate the effectiveness of the proposed filtering method on five datasets. We utilize $\rho_{\text{non-trans}}$, as defined in Equation 3, to quantify the degree of preference non-transitivity in the evaluation of LLMs. A higher value of $\rho_{\text{non-trans}}$ indicates a greater extent of non-transitive preferences. Additionally, we employ $\tau_{\text{avg}}$, as defined in Equation 8, to assess the overall clarity of preferences, where a lower value reflects improved preference clarity. Furthermore, we provide a comprehensive measure of the alignment between model preferences and human preferences by calculating the human agreement rate and the spearman correlation coefficient between the model’s evaluations and human assessments. Additionally, we investigate two scenarios in which Evaluator LLMs are prone to exhibit non-transitive preferences, see Appendix D.
@@ 5.1 Main Results
5.1.1 Analysis of Non-Transitivity and Clarity in Preferences
As shown in Table 1, across all testing sets, models fine-tuned on the "Cleaned" training set exhibit the lowest degree of preference non-transitivity and the highest clarity of preferences, surpassing even the teacher model Qwen2.5-Max. This highlights the effectiveness of our proposed data filtering method. Additionally, models fine-tuned on the "Raw" dataset perform comparably to the teacher model Qwen2.5-Max, demonstrating that data distillation can indeed enhance a model’s evaluation capability relative to its original performance. However, due to the inherent preference non-transitivity of the teacher model, direct distillation may learn low-quality preference data, thereby limiting the ability of the distilled model to surpass the teacher model and resulting in similar performance.
@@ 5.1.1 Analysis of Non-Transitivity and Clarity in Preferences
To validate that models fine-tuned on the "Cleaned" training set achieve lower preference non-transitivity and higher preference clarity even on "unseen" questions, we performed cross-validation by testing each fine-tuned model on five testing sets. As shown in Table 2, models fine-tuned on Cleaned data show a 13.78% average reduction in non-transitivity and a 0.0879 decrease in normalized structural entropy compared to those trained on Raw data, indicating clearer overall preferences. Detailed results are in Appendix B.3. Notably, even advanced Evaluator LLMs (e.g., Qwen2.5-Max at 67.96%) exhibit high non-transitivity and entropy (0.8591 and 0.8095, respectively), reflecting limited preference clarity.
Training Set $\rho_{\text{non-trans}}$ $\tau_{\text{avg}}$
Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
Helpful_Base 65.95% 50.33% -15.62% 0.8204 0.7173 -0.1031
Vicuna 61.56% 50.96% -10.60% 0.8063 0.7367 -0.0696
Oasst 62.61% 48.73% -13.88% 0.7966 0.7101 -0.0865
Koala 64.05% 48.65% -15.40% 0.8044 0.7177 -0.0867
Self-instruct 67.23% 53.84% -13.39% 0.8286 0.7351 -0.0935
Average 64.28% 50.50% -13.78% 0.8113 0.7234 -0.0879
Table 2: Comparison of average preference non-transitivity and average clarity between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct across testing sets. The leftmost column indicates the source dataset used for training, while other columns show model average performance on five testing sets. For example, the "Raw" column under Helpful_Base shows results for unfiltered data, and the "Cleaned" column reflects filtered data.
5.1.2 Human Agreement Rate and Spearman Correlation Analysis.
We evaluate the alignment of fine-tuned Evaluator LLMs with 2.5k human-annotated preference labels from the AlpacaEval dataset, which contains around 650 questions with two candidate responses each, annotated by four human evaluators Li et al. (2023). The questions are sourced from Helpful_Base, Oasst, Koala, Vicuna, Self-Instruct, and Alpaca.
Human Agreement Rate: This metric quantifies the alignment between Evaluator LLMs and human annotators’ majority preferences. For each instruction, four predictions are generated by the Evaluator LLM. Using a leave-one-out approach, each prediction is compared against the majority vote of the remaining three human annotations, excluding the corresponding human annotation. The final score is computed as the average accuracy across all four comparisons per instruction.
Spearman Correlation:This metric evaluates the Spearman correlation between rankings based on Evaluate LLMs’ preferences and those derived from human preferences.
As shown in Table 3, models fine-tuned on the corresponding "Cleaned" training sets of Vicuna, Oasst, koala and Self-instruct exhibit higher human agreement rate compared to those fine-tuned on the "Raw" training sets, with improvements ranging from 0.6% to 1.2%. Additionally, models fine-tuned on the corresponding "Cleaned" training sets of Helpful_Base, Vicuna, and Self-instruct exhibit higher Spearman correlation compared to those fine-tuned on the "Raw" training sets. Notably, on the Helpful_Base "Cleaned" training set achieves the largest improvement, with the Spearman correlation increasing from 0.93 to 0.97.
5.2 Ablation Experiments
Ablation of Different Base Models: We conducted repeated experiments using the LLaMA3.1-8B-Instruct model as the base model. As shown in Appendix B.2, models trained on the "Cleaned" training set outperform those trained on the "Raw" training set and also surpass the advanced open-source Evaluator LLM Auto-J 13B. We further validated that this superiority holds true even for "unseen" questions, with detailed results provided in Appendix B.3, further highlighting the importance of high-quality data in enhancing model performance.
Ablation of Different Data Filtering Methods: As shown in Figure 3, the "Cleaned" training set constitutes approximately 80% of the "Raw" training set. To validate the effectiveness of our data filtering method, we randomly filtered 20% of the training data and fine-tuned the Qwen2.5-7B-Instruct model using the remaining 80%, denoted as Qwen-Random. As shown in Table 1, the overall performance is similar to the model trained using the "Raw" training set, and showed higher preference non-transitivity on the Vicuna testing set. This comparison underscores the effectiveness of our targeted filtering strategy in enhancing data quality.
Ablation of Different Prompt Forms: We conducted additional experiments using a prompt format that includes ‘tie’ preference outcomes. The detailed prompt template is provided in Appendix C.2, and the experimental setup and results are presented in Appendix B.1, further validating the effectiveness of our data filtering method. Notably, when using prompts that include ‘tie’ preference outcomes, Evaluate LLMs exhibit higher preference non-transitivity and lower overall preference clarity compared to prompts without ‘tie’ outcomes. This finding is consistent with the results reported in prior work Xu et al. (2025).
Training Set Human Agreement Rate Spearman Correlation
Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
Helpful_Base 66.9% 66.9% +0.0% 0.93 0.97 +0.04
Vicuna 65.2% 65.8% +0.6% 0.97 0.98 +0.01
Oasst 66.4% 67.6% +1.2% 0.93 0.93 +0.00
Koala 66.4% 66.9% +0.5% 0.98 0.98 +0.00
Self-instruct 67.8% 68.3% +0.5% 0.98 1.00 +0.02
Average 66.5% 67.1% +0.6% 0.96 0.97 +0.01
@@ 5.2 Ablation Experiments
Table 3: Comparison of Human Agreement Rate and Spearman Correlation between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct. The structure matches Table 2.
@@ 6 Conclusion
In this paper, we present a detailed analysis of preference non-transitivity in pairwise evaluations by LLM-as-a-Judge using graph theory, while measuring overall clarity through the two-dimensional structural entropy of directed graphs. By filtering low-quality training data that induce non-transitive preferences, our method effectively reduces preference non-transitivity, enhances overall preference clarity, and improves alignment with human judgments. This principled framework contributes to more reliable automated evaluation with LLMs and offers valuable insights for future research on preference consistency and robust benchmarking.
Limitations and Future Work
Although we have validated the effectiveness of our filtering approach on five datasets from AlpacaEval, it still cannot fully cover all real-world scenarios, and further validation across a broader range of domains is needed. Additionally, while our method has successfully reduced the issue of preference intransitivity, it has not been completely eliminated. Further efforts to minimize preference intransitivity remain an important direction for future work. Lastly, our study primarily focused on pairwise comparison tasks in single-turn dialogues, exploring multi-turn dialogue tasks presents an intriguing avenue for future research.
Ethics Statement
This research focuses on analyzing and mitigating non-transitive preferences in LLMs when used as evaluators for open-ended tasks. The methodology involves processing publicly available data and conducting computational experiments using existing datasets and open-source frameworks, without involving human subjects or sensitive personal information. Therefore, no specific ethical approval was required for this study.
Appendix A LLM Details
This section provides a detailed overview of the models used to construct the training and testing sets in the experiment. We selected a total of 21 LLMs, covering advanced closed-source models, large-scale and small-scale open-source models, as well as dense and Mixture-of-Experts (MoE) architectures. Specifically, the selection includes the following:
  • OpenAI: OpenAI model series OpenAI et al. (2024).
  • Anthropic: Claude model seriesAnthropic (2023, 2024).
  • Mistral: Mistral model seriesJiang et al. (2024, 2023).
  • Alibaba: Qwen model seriesQwen Team (2024a, b).
  • Meta: LLaMA model series Meta AI (2024a, b); Touvron et al. (2023).
  • Yi: Yi model seriesAI et al. (2025).
  • Google: Gemini model seriesTeam et al. (2024).
  • Other Notable Models: Vicuna 33BChiang et al. (2023), Tulu 2+DPO 70BIvison et al. (2023), and WizardLM 13BXu et al. (2024a).
The pre-generated outputs for each model on the AlpacaEval dataset were sourced from the AlpacaEval leaderboard.
Appendix B Additional Experimental Results
B.1 Experimental Results of Different Prompt Forms.
Across five testing sets, models fine-tuned on the "Cleaned" training set exhibit the highest preference clarity. In four of these testing sets (excluding Vicuna), these models also demonstrate the lowest preference non-transitivity, outperforming models fine-tuned on the "Raw" training set and the teacher model Qwen2.5-Max. We observe that models fine-tuned on the "Raw" training set, as well as the teacher model Qwen2.5-Max, exhibit a higher rate of ties in preferences on the Vicuna testing set. The prevalence of such ties—often arising from SCCs formed by cases like $(A=B)$, $(B=C)$, and $(C=A)$—significantly reduces the degree of preference non-transitivity when these SCCs are removed. As shown in Table 4, we compare the probability of ties among three models on the Vicuna testing set. Furthermore, the $\tau_{\text{avg}}^{\downarrow}$ values of models fine-tuned on the "Cleaned" training set indicate a higher overall clarity in preferences.
Figure 4: Comparison of Data Volumes Between "Raw" and "Cleaned" Training Sets Across Different Datasets (Using the CoT Comparison (Tie Allowed) Prompt Template)
Model tie probability
Qwen2.5-Max 7.59%
Qwen-Raw 10.06%
Qwen-Cleaned 6.13%
Table 4: Comparison of the probability of ties for models on the Vicuna testing set.
B.2 Experimental Results of Different Base Model.
We repeated the experiment using the LLaMA3.1-8B-Instruct model as the base model. The test results are presented in Table 6. We also compared our results with the advanced open-source Evaluator LLM Auto-J 13B, which is fine-tuned based on the LLaMA-2-13B-Chat model.
Due to the limited context length of 2048 tokens supported by the advanced open-source evaluators PandaLM and JudgeLM, which is insufficient to cover the evaluation tasks in our testing set, we did not include them in our experiments.
B.3 "Unseen" question validation
Appendix C Evaluate Prompts
C.1 CoT Comparison
The CoT Comparison prompt template, identical to AlpacaEval 2.0 Li et al. (2023), is detailed in Table 9.
C.2 CoT Comparison (Tie Allowed)
The CoT Comparison (Tie Allowed) prompt template, as described in (Xu et al., 2025), is detailed in Table 10.
Appendix D Under What Scenarios Are Evaluator LLMs Prone to Non-Transitive Preferences?
Evaluator LLMs, much like human judges, tend to exhibit non-transitive preferences under certain conditions.
First, when the quality differences between responses fall below the Evaluator LLM’s just noticeable difference (JND) threshold (a concept widely used in psychophysics to describe the minimum difference perceptible to an observer Stern and Johnson (2010)) the model may "lose evaluative focus." Analogous to how humans struggle to differentiate between highly similar colors on a gradient from yellow to orange, the Evaluator LLM may perceive adjacent responses as "almost equally good." While individual comparisons appear consistent, cumulative evaluations can yield preference cycles (e.g., $A\succ B$, $B\succ C$, $C\succ A$), reflecting underlying non-transitivity.
Second, when responses differ significantly in stylistic dimensions, the Evaluator LLM may exhibit indecision due to the inherent conflict in multi-dimensional evaluation. Consider an open-ended question evaluated along three criteria: rigor, vividness, and richness:
  • Response A is the most rigorous, moderately vivid, and least rich.
  • Response B is the richest, moderately rigorous, and least vivid.
  • Response C is the most vivid, moderately rich, and least rigorous.
If overall preference is defined as outperforming another response in at least two out of three dimensions, instances such as $A\succ B$, $B\succ C$, and $C\succ A$ may arise—resembling the classic voting paradox, also known as the Condorcet paradox Fishburn (1974). This reflects a fundamental limitation in aggregating preferences across conflicting criteria: when each response possesses a distinct advantage, constructing a consistent global ranking becomes infeasible. The problem is especially pronounced in open-ended evaluations, where stylistic variance (e.g., concise and rigorous analysis vs. vivid and elaborate narratives) can lead to non-transitive outcomes in Evaluator LLM preferences.
Appendix E Algorithm Detail

### v3

@@ 3.2 Quality Analysis Framework for Evaluator LLM Training Data
Here, $\tau(G_{i})$ is the normalized structural entropy for the graph $G_{i}$ corresponding to question $q_{i}$, and $|\mathbf{Q}|$ is the total number of questions in the set $\mathbf{Q}$.
The $\tau_{\text{avg}}$ metric quantifies preference clarity: higher values indicate complex, non-transitive patterns, while lower values reflect relationships closer to linear order. It enables objective assessment of consistency across evaluator LLMs, where lower entropy denotes more coherent and interpretable preferences.
3.3 Filtering Strategy for Preference Data That Induce Non-Transitivity
If cycles within each SCC of a directed graph can be eliminated to form DAGs, the entire graph can be transformed into a DAG. Our framework leverages this property by converting each SCC into a DAG while preserving inter-SCC preference relations, ensuring global acyclicity without altering component relationships. From the resulting DAG, we derive transitive preference relations to filter evaluator LLMs’ preference data, enabling training data self-purification. The procedure is:
@@ 3.3 Filtering Strategy for Preference Data That Induce Non-Transitivity
    After processing all SCCs, the original cyclic graph is transformed into a DAG. This DAG is used to filter the training data as follows: For bidirectional edges $(v_{i}\leftrightarrow v_{j})$, the correct preferences are recorded as $\mathcal{J}(a_{i},a_{j})=\text{`tie'}$ and $\mathcal{J}(a_{j},a_{i})=\text{`tie'}$. For unidirectional edges $(v_{i}\rightarrow v_{j})$, the correct preferences are recorded as $\mathcal{J}(a_{i},a_{j})=\text{`lose'}$ and $\mathcal{J}(a_{j},a_{i})=\text{`win'}$.
@@
    The training data are traversed sequentially. All instances that are consistent with the correct preferences are added to the training set “Cleaned”, while the remaining data are placed in the training set “Discarded”.
The final “Cleaned” training data only retains linearly transitive preference data, resulting in $\rho_{\text{non-trans}}=0$ and $\tau_{\text{avg}}=0$. Our approach balances optimality and scalability, with an overall time complexity of $\mathcal{O}(|V|^{2}+|E|)$, consisting of $\mathcal{O}(|V|^{2})$ for graph construction, $\mathcal{O}(|V|+|E|)$ for SCC decomposition, $\mathcal{O}(|V|^{2})$ for intra-SCC reconstruction, and $\mathcal{O}(|E|)$ for edge filtering. The detailed algorithmic procedure is provided in Appendix A.
@@ 4 Experiment Setup
Qwen-Base 82.8% 0.922 78.9% 0.891 83.4% 0.914 81.0% 0.910 81.5% 0.912
Qwen-Raw 62.0% 0.796 57.5% 0.803 55.8% 0.773 64.3% 0.816 59.7% 0.767
Qwen-Random 60.6% 0.790 63.6% 0.808 58.8% 0.771 60.4% 0.812 57.1% 0.764
Qwen-Cleaned (ours) 44.9% 0.700 43.9% 0.726 47.0% 0.694 48.5% 0.715 49.0% 0.680
LLaMA-Base 76.4% 0.852 67.0% 0.846 67.4% 0.793 69.9% 0.818 71.0% 0.809
LLaMA-Raw 59.0% 0.765 60.2% 0.772 58.2% 0.760 57.2% 0.746 60.0% 0.783
LLaMA-Cleaned (ours) 40.2% 0.652 45.4% 0.691 43.0% 0.629 44.4% 0.659 42.0% 0.642
Table 1: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the “Raw”, “Random”, and “Cleaned” training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the ”Cleaned” training set derived from filtered Helpful_Base training set.
4.1 Dataset
In this study, we conduct experimental validation using the AlpacaEval benchmark (alpaca_eval). AlpacaEval is specifically designed to assess the overall capabilities of LLMs in open-ended tasks, covering a wide range of evaluation scenarios such as reasoning and text and code generation. The benchmark comprises five datasets, Helpful_Base, Oasst, Koala, Vicuna, and Self-Instruct.
4.2 Preference Data Collection
We evaluated Qwen2.5-Max (qwen25max) on 2.5k human-annotated samples from AlpacaEval using the chain-of-thought (COT) comparison (alpaca_eval). Qwen2.5-Max achieved a human agreement rate of 68.9%. Based on these results, we selected it as our teacher model for preference data generation.
4.3 Experiment Details
21 representative LLMs were selected from the AlpacaEval leaderboard (14 for training, 7 for testing). All experiments employed CoT comparison templates with temperature set to 0, and preferences were generated via the Qwen2.5-Max API. The unfiltered “Raw” datasets were processed according to the procedure described in Section 3.3 to produce “Cleaned” training sets; the corresponding distributions are illustrated in Figure 4. Qwen2.5-7B-Instruct was fine-tuned on both versions of each of the five datasets using LoRA (hu2022lora) (rank =8=8, 3 epochs, learning rate $1\times 10^{-4}$, batch size =16=16).
Figure 4: Comparison of data volumes between “Raw” and “Cleaned” training sets across datasets. The “Cleaned” training set’s volume is approximately 80% of the “Raw” training set for each dataset.
@@ 5 Results and Analysis
This section verifies the effectiveness of the proposed filtering method on five datasets. The analysis begins with an examination of preference non-transitivity degree and overall preference clarity of evaluator LLMs. To further demonstrate the method’s effectiveness, experiments based on actual usage scenarios are conducted using the MT-bench evaluation framework to calculate the standard deviation between final results. Additionally, a detailed data quality analysis examines two key aspects of manual evaluation: (1) consistency between manual annotators, and (2) consistency between evaluation results and human majority vote.
@@ 5.1 Main Results
Analysis of preference non-transitivity and clarity. As shown in Table 1, models fine-tuned on the “Cleaned” training set demonstrate the lowest preference non-transitivity and highest preference clarity across all datasets, including Qwen2.5-Max, highlighting the effectiveness of our proposed data filtering methodology. To verify that models fine-tuned on the “Cleaned” training set achieve lower preference non-transitivity and higher preference clarity even on “unseen” questions, we conducted cross-validation as our out-of-domain experiment by testing each fine-tuned model across five test sets. As demonstrated in Table 2, models fine-tuned on “Cleaned” training set exhibited an average reduction in non-transitivity of 13.8% compared to those trained on “Raw” training set. Table 3 shows a reduction of 0.088 in normalized structural entropy, indicating clearer overall preferences. Detailed results are in Appendix B.
Dataset Raw Cleaned $\Delta$
Helpful_Base 66.0% 50.3% -15.7%
Vicuna 61.6% 51.0% -10.6%
Oasst 62.6% 48.7% -13.9%
Koala 64.1% 48.7% -15.4%
Self-instruct 67.2% 53.8% -13.4%
Average 64.3% 50.5% -13.8%
@@
Table 2: Comparison of $\rho_{\text{non-trans}}$ (average preference non-transitivity) between models fine-tuned on “Raw” and “Cleaned” training sets using Qwen2.5-7B-Instruct.
Dataset Raw Cleaned $\Delta$
Helpful_Base 0.820 0.717 -0.103
Vicuna 0.806 0.737 -0.069
Oasst 0.797 0.710 -0.087
Koala 0.804 0.718 -0.086
Self-instruct 0.829 0.735 -0.094
Average 0.811 0.723 -0.088
Table 3: Comparison of $\tau_{\text{avg}}$ (average preference clarity) between models fine-tuned on “Raw” and “Cleaned” training sets using Qwen2.5-7B-Instruct.
Human validation of discarded data quality. To verify the low quality of discarded data, we evaluated 100 randomly sampled instances from both “Cleaned” and “Discarded” sets across five datasets (1,000 total instances). Three independent annotators assessed these samples, enabling measurement of inter-annotator consistency and model-human alignment. Results in Table 4 indicate that human annotator consistency in the “Discard” training set averages only 34.4%, significantly lower than the 52.6% observed in the “Cleaned” training set. This confirms that data leading to non-transitive preferences is inherently ambiguous. Furthermore, model evaluation reveals that consistency between the “Discarded” training set and human majority votes averages merely 51.2%, substantially below the 80.6% achieved by the “Cleaned” training set. These findings provide compelling evidence that the “Discarded” training set contains low-quality preference pairs unsuitable for training models to learn consensus human preferences.
For deeper insight into non-transitive preferences, we observed that source response pairs exhibiting non-transitivity demonstrate higher textual similarity. This observation suggests that preference cycles are likely to occur when the quality difference between source responses falls below the just noticeable difference (JND) (stern2010just) threshold of the evaluator LLM. As shown in Table 5, we quantified content similarity between response pairs by calculating Self-Bleu scores. Source response pairs with non-transitive preference relations consistently displayed higher Self-Bleu values than those with transitive relations, confirming that preference non-transitivity correlates with greater text similarity between comparative samples.
HC MHA
Dataset Cleaned Discarded Cleaned Discarded
Help. 59% 46% 72% 45%
Vicuna 48% 24% 78% 47%
Oasst 56% 33% 83% 53%
Koala 54% 41% 84% 61%
Self. 46% 28% 86% 50%
Average 52.6% 34.4% 80.6% 51.2%
Table 4: Comparison of Human Evaluation Consistency and Model-Human Agreement between training sets. HC: Human Consistency (inter-annotator consensus); MHA: Model-Human Agreement (alignment with majority vote). Help. = Helpful_Base; Self. = Self-instruct.
Pref. Helpful. Vicuna Oasst Koala Self.
Non. 0.0903 0.1026 0.0946 0.0888 0.1135
Trans. 0.0713 0.0918 0.0788 0.0746 0.0829
Table 5: Comparison of source text pair Self-Bleu scores across datasets. Non. and Trans. indicate source pairs from non-transitive and transitive preferences; Helpful. = Helpful_Base; Self. = Self-instruct.
Performance analysis of real-world evaluation systems. We compared the performance of our trained evaluator LLMs within the MT-bench evaluation framework (zheng2024judging_mt). We adopted the adjusted win rate calculation consistent with the MT-bench evaluation methodology. This metric normalizes performance by treating ties as partial wins, defined as $r_{\text{adj}}=(r_{w}+0.5\cdot r_{t})/(r_{w}+r_{l}+r_{t})$, where $r_{w}$, $r_{l}$, and $r_{t}$ represent the win rate, loss rate, and tie rate respectively. The adjusted rate $r_{\text{adj}}$ serves as the primary indicator for our model ranking. Results in Table 6 demonstrate that models fine-tuned on the “Cleaned” training sets exhibit significantly higher standard deviations in their adjusted win rates across five datasets. This indicates enhanced robustness of the evaluation results and better differentiation of performance differences between models.
Dataset Help. Vicu. Oasst Koala Self.
Cleaned 16.1% 16.6% 13.1% 14.9% 9.7%
Raw 12.2% 13.6% 11.0% 12.3% 8.2%
$\Delta$ +3.9% +3.0% +2.1% +2.6% +1.5%
Table 6: Standard deviation analysis. Help. = Helpful_Base; Vicu. = Vicuna; Self. = Self-instruct.
MHA SC
Dataset Raw Cleaned Raw Cleaned
Helpful_base 66.9% 66.9% 0.93 0.97
Vicuna 65.2% 65.8% 0.97 0.98
Oasst 66.4% 67.6% 0.93 0.93
Koala 66.4% 66.9% 0.98 0.98
Self-instruct 67.8% 68.3% 0.98 1.00
Average 66.5% 67.1% 0.96 0.97
Table 7: Comparison of Model-Human Agreement (MHA) and Spearman Correlation (SC) between fine-tuned models. Spearman Correlation quantifies the rank correlation between model and human preferences.
Analysis of human agreement impact. The model’s alignment with human preferences was evaluated by calculating agreement rates and Spearman correlation coefficients on 2.5k manually annotated preference labels from AlpacaEval. Results in Table 7 indicate that models fine-tuned on the “Cleaned” training set consistently outperformed those trained on “Raw” training set across multiple dataset, with maximum improvements of 1.2% in human agreement and 0.04 in Spearman correlation. These results demonstrate that filtering ambiguous, low-quality preference data not only improves discrimination and enhances the performance of downstream evaluation systems but also increases consistency with human judgment, providing strong evidence for practical applications.
5.2 Ablation Studies
Effect of Different Base Models: Testing using the same steps with LLaMA3.1-8B-Instruct confirms our findings: models trained on the “Cleaned” training set consistently outperformed those trained on the “Raw” training set (Table 1). Data Filtering Analysis: Our “Cleaned” set retained approximately 80% of the raw data (Figure 4). A control experiment with random 20% filtering (Qwen-Random) produced results similar to training on the “Raw” set but with higher preference non-transitivity, confirming that our targeted approach improves data quality beyond simple reduction (Table 1). Prompt Format Variations: Additional experiments with prompts explicitly allowing “tie” judgments further validated our methodology. Detailed prompt templates and results are provided in Appendices B and C.
@@ 6 Conclusion
We propose ELSPR, a graph-theoretic framework to analyze and mitigate non-transitivity in evaluator LLMs by modeling pairwise comparisons as a tournament graph and filtering problematic data. Ambiguous, low-quality preferences are a major source of non-transitive judgments, and targeted filtering improves evaluation reliability. Empirical results show that filtered data aligns better with human judgments and can fine-tune models to reduce non-transitive bias and structural entropy. This work underscores the importance of data quality and provides a scalable approach to enhance consistency on future benchmarks.
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant Nos. 62572105 and U22B2005; the Liaoning Revitalization Talents Program under Grant No. XLYC2403086.
Appendix A Algorithm Detail
In this section, we provide the detailed pseudocode for the filtering algorithm described in Section 3.3 of the main text.